OpenStudy (anonymous):

What is the value of a(b - c) if a = 2, b = -3 and c = -2?

jimthompson5910 (jim_thompson5910):

Replace every copy of 'a' with 2 Replace every copy of b with -3 Replace every copy of c with -2 Then evaluate a(b - c) = 2*(-3 - (-2)) = ??

OpenStudy (anonymous):

-2